### v1

# Democratizing Signal Processing and Machine Learning: Math Learning Equity for Elementary and Middle School Students

Abstract
Signal Processing (SP) and Machine Learning (ML) rely on good math and coding knowledge, in particular, linear algebra, probability, and complex numbers. A good grasp of these relies on scalar algebra learned in middle school. The ability to understand and use scalar algebra well, in turn, relies on a good foundation in basic arithmetic. Because of various systemic barriers, many students are not able to build a strong foundation in arithmetic in elementary school. This leads them to struggle with algebra and everything after that. Since math learning is cumulative, the gap between those without a strong early foundation and everyone else keeps increasing over the school years and becomes difficult to fill in college. In this article we discuss how SP faculty and graduate students can play an important role in starting, and participating in, university-run (or other) out-of-school math support programs to supplement students’ learning. Two example programs run by the authors (CyMath at ISU and Ab7G at Purdue) are briefly described. The second goal of this article is to use our perspective as SP, and engineering, educators who have seen the long-term impact of elementary school math teaching policies, to provide some simple almost zero cost suggestions that elementary schools could adopt to improve math learning: (i) more math practice in school, (ii) send small amounts of homework (individual work is critical in math), and (iii) parent awareness (math resources, need for early math foundation, clear in-school test information and sharing of feedback from the tests). In summary, good early math support (in school and through out-of-school programs) can help make SP and ML more accessible.
Submitted to IEEE Signal Processing Magazine for consideration as a Perspectives article (within In the Spotlight) or as an SP Education article (within Learning SP).
About the authors. Vaswani is a Professor of Electrical and Computer Engineering (ECE) at Iowa State University, where her teaching and research lie in statistical signal processing (SP) and machine learning (ML). She is a Fellow of IEEE for contributions to dynamic structured high-dimensional data recovery. In the last 20 years, she has extensively published in various IEEE Transactions and has taught multiple Math for SP and ML courses. These include undergraduate Probability, an undergraduate course on ML from an SP perspective, graduate level Detection and Estimation Theory, and High-dimensional Probability for ML and SP. In 2020, she founded and directs and runs the CyMath (graduate student led) elementary and middle school mathematics (math) tutoring and mentoring program. Selim is a networking and computer engineering educator and researcher and an Associate Teaching Professor at Iowa State since 2018. He has been a part of CyMath since 2024 and now leads its online tutoring component. Gibert is a Program Manager for the Minority Engineering Program at Purdue University and a former middle school mathematics teacher. She coordinates the Algebra by Seventh Grade (Ab7G) program at Purdue, which is a much larger scale math mentoring program for school students.
Introduction
This article discusses how early interventions in math education can bridge learning gaps and promote equity in signal processing (SP) and machine learning (ML). We first explain why early math is critical for SP and ML in Sec. I. In Sec. II, we describe (well-known and lesser known) systemic barriers that prevent some students from getting a good early math foundation. We then describe how SP faculty, graduate students can play an important role in starting, and participating in, university-run (or other) out-of-school math support programs to supplement students’ learning (Sec. III). Two example programs are briefly described also. In Sec. IV, we use our perspective as college educators who have seen the long-term impact of current elementary math education to provide some simple zero-cost suggestions for public schools that may help reduce some of the learning gaps. The first section below can be appropriately modified or ignored and the rest of the article will still be relevant for most science, technology, engineering, and mathematics (STEM) fields.
I Why Early Math Education matters for Signal Processing (SP) and Machine Learning (ML)
Understanding modern signal processing (SP) and machine learning (ML) – basic SP, statistical SP, digital SP, estimation and detection theory, or ML algorithms – requires a good grasp of linear algebra (vector and matrix algebra), probability and statistics, and complex numbers, and an ability to code algorithms based on these mathematics (math) concepts in MATlAB or Python. The latter is essential to understand, develop, and evaluate novel applications of SP and ML algorithms. However, one cannot learn linear algebra, or even learn to code it in, without a strong knowledge of basic scalar algebra that is typically taught in middle school, and sometimes in high school. Algebra also forms the foundation of the ability to understand abstract probability concepts such as probability mass function or probability density function. Algebra, in turn, cannot easily make sense to a student without fluency in elementary school arithmetic (add, subtract, multiply divide; negative numbers; fractions, decimals, and four operations with these; exponents). As an example, the ability to solve for $x,y$ from two equations in $x,y$: 0.2($30.2(3x+5)-9x+22.1=1/$ 3 and $x-y=0.29$ requires arithmetic fluency. Solving for two variables from two linear equations forms the basis of understanding linear algebra, e.g., how to solve for an unknown vector $\bf{x}$ from $\bf{A}\bf{x}=\bf{b}$. This, in turn, is the basis for understanding Least Squares estimation which is used for parameter learning for Linear Regression. This understanding is essential to correctly code in the least squares estimation algorithm, or to correctly use software tools (e.g., the backslash operator in MATLAB or operators from the NumPy package in Python) that provide such a solution. In fact, in the ChatGPT era, where code snippets and code syntax can be automatically generated, math will become even more important because one needs the skills to verify whether the auto-generated code is indeed mathematically correct.
As an example, implementing least squares estimation to solve for $\bf{x}$ from $\bf{b}:=\bf{A}\bf{x}+\bf{w}$, where $\bf{w}$ is modeling error and $\bf{A}$ is a tall matrix, requires computing $\bf{A}^{\dagger}\bf{b}:=(\bf{A}^{\top}\bf{A})^{-1}\bf{A}^{\top}\bf{b}$. This can be coded in various ways in Python, but some are better than others in certain settings. For instance, the pinv function from some of the Python packages does not always work as desired. Moreover, the condition number of $\bf{A}$ governs solution accuracy, and when it is ill-conditioned, none of the approaches give an accurate answer. “Regularization” approaches exist to deal with ill-conditioning. Realizing all this requires a good understanding of linear algebra.
In summary, SP and ML rely on good college-level math knowledge. Since math learning is cumulative, the gap between those who did not get an opportunity to learn arithmetic, and hence scalar algebra, well and everyone else keeps increasing over the school years [1] and becomes difficult to fill by the time they reach college. We provide a visual in Fig. 1 that shows this. One way to make SP learning more accessible is to fill the math learning gap while it is still small by providing extra support early.
Figure 1: Figure indicating that math learning is cumulative. This is taken from https://www.educationnext.org/grade-level-expectations-trap-how-lockstep-math-lessons-leave-students-behind/ which uses data from [1].
II Systemic Barriers
Due to various systemic barriers encountered by students from under-represented backgrounds, many do not get a good early foundation in mathematics and consequently, opportunities for economic mobility via higher education are blocked to them [2]. Well-known barriers include inadequate resources for public education, family income and education levels, non-uniform teacher quality, COVID-19 learning losses, unconscious bias, language barriers and a culture of low expectations [3]. In addition to the above, the authors of this article have also noticed the following lesser-known barriers, which should be easier to remove since they are due to design changes that were introduced to improve the elementary school experience. These include: (i) insufficient focus on math practice in elementary school (both during school hours or as homework); (ii) little information provided to families that a good grasp of elementary school math is essential for later success, how to help students get this grasp, or that students are often tested in fourth or fifth grade in order to allow them to join advanced math tracks starting in middle school in many countries. (iii) Moreover in some districts this track has been removed and this can be even worse for the high-performing students in these schools. This track, that allows students to take algebra by 7th or 8th grade, is an unsaid requirement for students to be able to do the advanced math courses in high school that are needed for success in SP, ML and, in fact all science, technology, engineering and math (STEM) fields. Access to algebra has been validly referred to as the greatest civil rights issue of our time as it allows students access to economic ladders of opportunities [2]. The last issue may be US-specific.
In the next two sections below, we discuss two types of approaches that may reduce the impact of some of these barriers.
  1. 1.
    One is via out-of-school programs that SP majors (undergraduate or graduate students, faculty, or industry professionals), and anyone with college-level math training and passion for math, can start or participate in. University support can make it easier to start such a program. We describe two such programs in Sec. III next.
  2. 2.
    However, no out-of-school program can achieve the broad reach and accessibility of public schools in serving all students. In Sec IV, we share our suggestions for public schools. As signal processing and engineering educators, who have been running school math support programs, these suggestions are based on what we see as the long-term impact of elementary school math teaching policies.
III Partial solution 1: out-of-school math programs such as CyMath and Ab7G
We describe two out-of-school university run math programs – Algebra by 7th Grade (Ab7G) and CyMath – here. Ab7G is an elementary and middle school math mentoring program that was started at Purdue in 2017 by their Minority Engineering Program (MEP) and is coordinated by Gibert. Most of the mentors are undergraduate students. CyMath is Iowa State University’s school math tutoring program that was started by Vaswani in Fall 2020 and that primarily relies on volunteer graduate students or faculty as the tutors. It has been expanded during 2023-24 with help from Purdue’s Ab7G program and Gibert. Selim joined CyMath in 2024 and now leads its online tutoring component. It is a smaller program and provides twice-a-week in-person or hybrid support, while Ab7G is a much larger program that provides monthly support. Both provide free regular math support, along with at-home learning resources, to students in grades 2-8 while prioritizing those from backgrounds under-represented in Engineering and Computing.
Algebra by 7th Grade (Ab7G) - Purdue. Ab7G was established at Purdue in 2017. Chevron and Duke Energy seeded the funding for Ab7G with a generous grant. The mission of Ab7G is to increase the number of under-represented 7th grade students in the United States that are academically prepared to take algebra. The program is provided at no cost to participants. Under the leadership of Director Virginia Booth Womack of the Minority Engineering Program (MEP) in 2017, forty-four students in grades three through seven from local school districts near Purdue’s West Lafayette campus participated as the inaugural cohort. A partnership was created with Lafayette School Corporation (LSC) to help recruit students for the program. Students attended in-person sessions twice per month for a total of sixteen sessions during the 2017-2018 academic calendar. In 2024, Ab7G had 208 participants, including some virtual-only participants from outside Indiana and outside the US also.
The exigencies of the COVID-19 pandemic forced the program to move to a virtual platform. Currently, the program operates in a hybrid format, offering two in-person sessions each semester, accompanied by an asynchronous virtual option during the in-person sessions. Additionally, four virtual sessions are provided each semester–fall and spring. The Saturday sessions are structured for students to work on McGraw Hill’s ALEKS (Assessment and Learning in Knowledge Spaces) adaptive learning mathematics software, mentorship with an undergraduate student, in-person sessions lunch, and a STEM activity. Ab7G participants are encouraged to work at least fifteen minutes per day most days a week on ALEKS outside of the Saturday program and attempt to master three topics per week. At the beginning of each session, students are recognized for reaching their time and topic goals— honorable mentions are given to students that come close. The program is built upon three foundational pillars: student self-efficacy, mentorship, and parental engagement. Students enhance their mathematical proficiency through both online and in-person instruction, collaborative activities, practical projects, and opportunities to interact with qualified STEM experts. Parent workshops involve exploring mathematical principles, navigating the online math tool, completing practical projects specifically created to enhance their child’s learning, developing awareness for adolescent mental health, and creating growth mindsets.
Womack et al [4] reported significant growth in mathematics proficiency among students who completed a full grade level in ALEKS each year, n=20. This growth demonstrated a direct correlation with achieving proficiency or higher on Indiana’s Learning Evaluation and Assessment Readiness Network (ILEARN) test which is administered to Indiana students in grades 3-8 at the end of each school year.
CyMath - Iowa State. CyMath was founded in Fall 2020 to help support math learning for students from disadvantaged backgrounds during COVID-19. Most CyMath tutors are STEM graduate student volunteers. It started in online mode for one Des Moines, Iowa school in 2020 and ran that way until 2022. Attendance was excellent in the first year but dropped in the second year. In 2023 Fall, with using advice from Ab7G, Vaswani re-started CyMath as an in-person program for the local Ames Iowa school district. The goal of CyMath is to begin tutoring students in third or fourth grade and follow them through the school years, at least until the end middle school. Its long-term goal is to increase the number of youth prepared to pursue, and thrive in, math-intensive majors in college.
In the online mode (2020-2022), tutoring was provided to groups of two or three students once a week. In the in-person mode (2023-24 school year), tutoring was fully in-person and one-on-one. Each student got one session a week in Fall 2023, and two a week in Spring 2024. Tutoring was followed by a 20-minute play time with LEGOs or on the school playground. Tutoring ran after-school at the school itself and busing was provided by the district to take students back home, thus reducing demands on working parent(s) time. CyMath benefited significantly from the school principal’s support. Zoom sessions, and some in-person ones (at Iowa State), were also run over Summer 2024. We taught a total of about 25 students for some part of the 2023-24 school year. New students were added as soon as more tutors became available or when some students left (September, January, March, May). For the 2024-25 school year, CyMath is transitioning to hybrid mode tutoring. Each week, one session will be in-person and one online. Hybrid tutoring helps with reaching parents better, maintaining tutor continuity after the tutor graduates, and should help CyMath expand more easily to schools that are farther away.
CyMath tutors are volunteer graduate students, and some faculty, from Electrical and Computer Engineering (signal processing primarily), Mathematics, Statistics. These are supported by a few (paid) undergraduate students from Elementary Education, and related, majors who help students (children) stay focused, and also serve as tutors where needed. Regular text messages and phone calls, along with email, are used to motivate and encourage parents to help their students also do at-home math practice. Borrowing from Ab7G, CyMath buys accounts for the adaptive math learning software ALEKS www.aleks.com. Tutors use this as a base for tutoring, along with also encouraging students to regularly work on math on ALEKS at home. Roughly 20 minutes a day for about five days a week is recommended. Many students do not have reliable internet access and hence tutors also send home workbooks and worksheets (printed from websites such as k5learning.com). CyMath does not insist on a particular tutoring curriculum, or use of ALEKS, and so far has provided very limited training to tutors. It relies on the assumption that tutors know their math (have graduate level training), have a passion for it, and are responsible adults (tutors are carefully chosen and observed). CyMath also informs parents and the students about the dates for state-wide testing and explains how the test runs (what adaptive testing means for example). A large fraction of tutors are international graduate students or immigrant faculty (including Vaswani and Selim). Since the language of math is universal, these tutors have been as effective as domestic graduate students. In fact, they also bring an international perspective to math learning and expose students to world cultures. At the same time, the tutoring helps the tutors become better college educators (teaching assistants or instructors).
As of May 2024, a total of eight third or fourth grade students had spent at least a semester at CyMath, and six of these had spent a full year. We looked at the MAP (Measures of Academic Progress) test score data for these eight students. MAP is a standardized test given to all Iowa students three times a year. Looking at two rounds of MAP test score data (January 2024 and May 2024), and comparing them against September 2023 (before CyMath started), we observed the following. Of the six full year students, one showed a 30 point growth, while the average and median for the full year for the school was 17 points. Two others showed 21 and 18 points growth. Of the two one semester students, one showed significantly above average growth.
Other Math tutoring projects. A math tutoring project was run by Anubha Gupta, Professor and SP researcher at IIIT-Delhi in India during 2022. Undergraduate student volunteers tutored students from a government school that is near IIIT-Delhi.
With writing this article our hope is to learn about other math support programs that SP and ML colleagues around the world participate in or run. If universities and colleges can partner with school districts and provide logistical support to their faculty, starting and running such programs can be a lot easier. Alternatively, schools, after-school child care programs, public libraries, institutions of religious worship, or non-profits could start similar programs and reach out to nearby colleges and universities to find tutors and mentors.
IV Partial solution-2: suggestions for schools based on lessons learned from CyMath, Ab7G, and engineering classrooms
As noted earlier, out-of-school programs cannot reach all the students that public schools reach. Even if there were infinite resources available to teach and transport all children, many will not want to, or will not be able to, join one due to other commitments such as having to work or to care for little siblings. In this section, we share our perspective as SP and engineering educators who have also been running school math support programs.Our perspective relies on the following facts.
(1) All school staff – teachers, principals, and higher-level administrators – want to help their students and improve their learning. In attempting to find better ways to improve things, they rely on recent elementary education literature. However, this literature largely consists of short-term studies on improving one aspect of learning, without regard to others. The reason is it is not practical to do long-term research studies since all students, especially those from disadvantaged backgrounds, move schools often; school systems are decentralized; and school districts often change curriculum every few years. This literature thus misses the long-term (elementary school teaching influencing college outcomes) effect of its recommended policies. As an example, [5] insists that homework in elementary school is not useful by looking at test results from just one elementary school grade. Different other studies argue that homework goes against equity because some students can get help at home while others cannot. However, it misses the point that many parents will make their child practice math at home even without homework.
(2) All students can be good enough in math. Some will be better than others, just as some are better at music or sports than others. This is acceptable as long as everyone gets sufficient and equitable chances to do their personal best. All should get a chance to show their real potential. In fact, “equity” means that those with disadvantage should be given extra support. For math learning equity, the support is likely to be much more effective if it is provided to students starting in younger grades.
(3) All families want the best for their child. One cannot prevent those with more resources – more time, more math knowledge, more money, or more awareness and knowledge of the school system –- to make their child learn more at home, or to hire paid tutors, or to get the school teachers to provide the necessary help to their child. While excessive parental intervention may be counter-productive, up to a point, this extra support benefits the child’s learning. This is especially the case if the child’s elementary school is teaching much less than what is common around the world at similar grade levels.
(4) College curriculums are similar across the world. College expectations (doing homework or studying before tests) are also similar. To a large extent, the same is true for the advanced tracks within secondary school math curriculums too. Consequently, a huge gap is created for students if their elementary schools do not sufficiently emphasize math, and if they do not get extra math support outside of school.
Some of the math knowledge gaps can be reduced with simple modifications to how many elementary schools operate. Most of our suggestions do not require extra staff and cost almost nothing. As signal processing and engineering educators, who also run school math support programs, we are providing suggestions based on how we see elementary school math learning influencing college success in signal processing and in all math-based course-work.
  • Practice in school. Our first suggestion is that elementary schools make students do enough math practice during school itself. This would be easy for teachers to incorporate if they do not have to convert every math lesson into a game. Planning for fun math games for each topic takes a lot of the teacher’s time and uses up a lot of classroom time as well. It likely takes away the class time that could otherwise be spent by the students on math practice; and by the teachers on checking the students’ work and correcting math mistakes early. In fact, reducing the time and effort needed to make all learning fun may also leave the teacher with more time for one-on-one support for students who need that.
  • Homework. Our second suggestion is that schools send reasonable amounts of math work home and require students to complete it. Individual math practice is essential for math fluency. Moreover, teachers should be allowed time to check students’ work carefully so that student learning can be corrected early.
  • Awareness. Most parents want the information needed to help their child learn and do better in school and later life. However, many parents are unaware that math learning is cumulative and hence a good early foundation in math is important, or that there are two or more math tracks starting in fifth or sixth grade in many countries. They also do not know how to help their child learn math better. This knowledge should not be the privilege of STEM professional parents or their relatives. It should be provided to everyone by the schools. Moreover, it will automatically get conveyed if math is emphasised, homework is sent, and free workbooks, printed worksheets, and information about good online resources such as Khan academy, are shared with families. These cost almost nothing. Math should be the least expensive subject to improve outcomes in, since it requires no equipment or supplies. Some information is summarized here Math-info.
  • Math tests awareness and feedback. This last suggestion may apply to the US only. There are few regular low-stakes tests in elementary schools, and, even when there are, very little feedback is provided to parents based on the test results. At the same time, schools do use state-wide tests, such as the Measure of Academic Performance (MAP) test, given a few times a year to evaluate students and to place them in different math tracks. However parents are not made sufficiently aware of this fact ahead of time. Students, and parents, are often not sufficiently informed about the test dates or that the tests are adaptive (and what that means). This is done in order to not put any demands on parents’ time, and follows advice provided by published literature. At the same time, if a parent checks with the teacher, they are provided all the information. Consequently, we end up disadvantaging a much larger number of children who may have otherwise come better prepared, or at least better rested, for these tests. Good test taking skills, and studying for a test, are also important for college success.
  • Use of summer to make-up for lost learning; relying on STEM students and professionals. Despite all efforts, some students may require additional support to stay on track. There should also be easily available literature, and resources, for helping students catch-up with lost learning at various stages. For example, for a pre-algebra student with learning gaps, as long as they can learn at least one way of doing all the arithmetic, middle school algebra will make sense to them. The summer before middle school can be used efficiently to help many students catch up. Summer is also the time many undergraduate and graduate students and faculty are more available and schools can tap into these resources by partnering with nearly colleges or universities.